EduFarming
Your list · 0हिं

Type at least two letters. Press Esc to close.

↑↓ navigate↵ openTAB cycle filter

Conditional Probability & Bayes' Theorem · Question 3 of 125

Question 3

EasyOriginal question

P(A)=713P(A) = \frac{7}{13}, P(B)=913P(B) = \frac{9}{13} and P(A∩B)=413P(A \cap B) = \frac{4}{13}. What is P(B∣A)P(B \mid A)?

Options for question 3

Show the answer and explanation
CorrectNot quiteAnswer C

Correct answer: C

Explanation

  1. P(A∣B)=P(A∩B)P(B)P(A \mid B) = \frac{P(A \cap B)}{P(B)}, and P(A′∣B)=1−P(A∣B)P(A' \mid B) = 1 - P(A \mid B).

  2. P(B∣A)=413713=47P(B \mid A) = \frac{\frac{4}{13}}{\frac{7}{13}} = \frac{4}{7}.

Worth noting

49\frac{4}{9} is P(A∣B)P(A \mid B); given AA, divide by P(A)P(A).

Concept
Conditional probability from given probabilities
Typical time
60 s
Your time
—
The rule this testsConditional probability
Practise this chapter →

Ten questions from Conditional Probability & Bayes' Theorem, at your own pace.

Practise this chapter →

Ten questions from Conditional Probability & Bayes' Theorem, at your own pace.